A mixture of 8% disinfectant solution is to be made from 10% and 5% disinfectant solution. How much of each solution should be used if 25 gallons if 8% solution are needed?
Answer by Alan3354(69443)

A mixture of 8% disinfectant solution is to be made from 10% and 5% disinfectant solution. How much of each solution should be used if 25 gallons if 8% solution are needed?
----------------
t + f = 25 --- total solution
10t + 5f = 8*25 --- concentration